OpenStudy (anonymous):

how is it possible to get point slope form out of (-4,2) (-1,7)

OpenStudy (anonymous):

y2-y2/x2-x1 so place in values 7-2 -1--4= 5/3

OpenStudy (mandre):

You know waht point slope form is?

OpenStudy (anonymous):

face plam use slope as m and your points pluged in y-y1= m(x-x1)